OpenStudy (anonymous):

Solve the following simultaneous equations: 3x-2y=y(x+2y)=4 @dan815 @Kainui @perl

OpenStudy (perl):

this is one equation 3x-2y=y(x+2y)=4

OpenStudy (perl):

did you mean to write 3x-2y = y x + 2y = 4 ?

OpenStudy (anonymous):

The question written 3x-2y=y(x+2y)=4 @perl

OpenStudy (anonymous):

@perl